# A Novel Technosignature Search in the Breakthrough Listen Green Bank Telescope Archive

### v1

@@ 5.6 Pipeline Limitations
While we detect many candidates that appear promising, we acknowledge several limitations in our current pipeline. We also suggest potential ways to address these limitations, although leave the implementation of these approaches to future work.
Our algorithm cannot find ET signals if they occur in especially dense RFI regions. Specifically, if there is persistent RFI within $\sim 1.5$ kHz of a promising signal, the frequency block with the promising candidate will not pass from Level 0 to Level 1, as the kurtosis of the OFF observations will be large due to the RFI. In these cases, a more sophisticated approach to comparing the intensity distributions of the ON and OFF scans could be developed. One potential method would be to compute the Pearson cross-correlations between the ON scans and also between the OFF scans, and highlight frequency blocks with high correlations between all of their ON scans but low correlation between their ON and OFF scans. Another way to address this problem would be to run successive iterations of the pipeline using varying block sizes. Smaller block sizes would make it more sensitive to signals in RFI dense regions, provided they are not drifting too strongly.
Another limitation is that we have excluded a signficant portion of the GBT’s frequency bandwidth from analysis. Only 70.5% of the total frequency coverage from $1.1-11.2$ GHz is passed into our pipeline. We remove these high-RFI regions as our pipeline performs poorly in especially dense RFI regions (as noted above), and because this $30\%$ of the total frequency coverage accounts for $\sim 95\%$ of all signals which would be computationally prohibitive to analyze. These regions can easily be included for analysis, though at the expense of compute time, which may be a worthwhile trade-off for smaller surveys.
A third limitation is bias towards certain types of signals. While our kurtosis criteria do not strongly preference any specific signal morphology, our additional vetting filters do to some extent. Our broadband filter tends to select signals that do not vary dramatically over time. While this is necessary to reduce the large number of broadband false positives, it could also reject promising signals that are intermittent. Despite this, our algorithm finds a wider variety of signal morphologies than previous SETI searches focused on narrowband Doppler-drifting signals, while still detecting many such signals. Our pipeline is also less likely than turboSETIto find “candidate” events that clearly contain persistent signals in both ON and OFF scans on visual inspection of spectrograms.
@@
While we use kurtosis as a metric, there are multiple other statistical measures that can be used to evaluate whether a spectrogram differs significantly from Gaussian noise. As one example, the Kullback–Leibler divergence could be calculated instead to determine the deviation of a frequency block’s intensity distribution from a Gaussian distribution. We leave the implementation of these other approaches to future work.
@@ 6 Conclusion
As part of the Breakthrough Listen program, we searched 3,077 nearby stars, using data from the GBT. We used four receivers, spanning a combined range of 1.10–11.2 GHz, and found no compelling candidates that are not attributable to RFI. We present and implement a novel pipeline, designed to search for signals present in the ON scans of observation cadences and not in the OFF scans. This is a search method less biased towards any specific signal type than many previous SETI searches. We analyze 9,684 unique cadences, representing 4,477 hours of on-sky observation time. We identify 11,913 frequency blocks 2.86 kHz wide deemed likely to contain a promising candidate signal. We visually inspected all of these candidates, and determined that all can likely be attributed to RFI. This work highlights the wide variety of RFI that hinders radio SETI searches, and we hope the techniques developed may prove useful in future studies dealing with increasingly heavy RFI environments.
This study constitutes the most comprehensive survey yet for radio evidence of advanced life around nearby stars. We improve on the results of Price et al. (2020) in terms of sensitivity and number of stars. Our novel search approach also makes us sensitive to a larger range of promising signals than previous surveys. The Drake Figure of Merit of our search is $8.4\times 10^{8}$, the Continuous Waveform Transmittter Rate Figures of Merit of our bands fall between $0.25-0.45$, and we probe down to EIRPs of $2\times 10^{12}$ W for a large numbers of stars ($\sim 1000$ stars at each band). With respect to each of these metrics, this SETI survey is the most sensitive and thorough to date. We add our pipeline to the many algorithms that have been developed to detect technosignatures in radio observations, and hope that this study encourages ensuing SETI surveys to take a more general approach to recovering potential ET signals from complex datasets.
@@ 7 Acknowledgements
The Breakthrough Prize Foundation funds the Breakthrough Initiatives, which manages Breakthrough Listen. The Green Bank Observatory facility is supported by the National Science Foundation, and is operated by Associated Universities, Inc., under a cooperative agreement. CP was funded as a participant in the Berkeley SETI Research Center Research Experience for Undergraduates Site, supported by the National Science Foundation under Grant No. 2244242.
8 Appendix
8.1 Quality Candidates

### v2

@@ 5.6 Pipeline Limitations
While we detect many candidates that appear interesting upon visual inspection, we acknowledge several limitations in our current pipeline. We also suggest potential ways to address these limitations, although leave the implementation of these approaches to future work.
Our algorithm cannot find ET signals if they occur in especially dense RFI regions. Specifically, if there is persistent RFI within $\sim 1.5$ kHz of a candidate signal, the frequency block with the candidate will not pass from Level 0 to Level 1, as the kurtosis of the OFF observations will be large due to the RFI. In these cases, a more sophisticated approach to comparing the intensity distributions of the ON and OFF scans could be developed. One potential method would be to compute the Pearson cross-correlations between the ON scans and also between the OFF scans, and highlight frequency blocks with high correlations between all of their ON scans but low correlation between their ON and OFF scans. Another way to address this problem would be to run successive iterations of the pipeline using varying block sizes. Smaller block sizes would make it more sensitive to signals in RFI dense regions, provided they are not drifting too strongly.
Another limitation is that we have excluded a signficant portion of the GBT’s frequency bandwidth from analysis. Only 70.5% of the total frequency coverage from 1.111.21.1-11.2 GHz is passed into our pipeline. We remove these high-RFI regions as our pipeline performs poorly in especially dense RFI regions (as noted above), and because this $30\%$ of the total frequency coverage accounts for $\sim 95\%$ of all signals which would be computationally prohibitive to analyze. These regions can easily be included for analysis, though at the expense of compute time, which may be a worthwhile trade-off for smaller surveys.
A third limitation is bias towards certain types of signals. While our kurtosis criteria do not strongly preference any specific signal morphology, our additional vetting filters do to some extent. Our broadband filter tends to select signals that do not vary dramatically over time. While this is necessary to reduce the large number of broadband false positives, it could also reject genuine signals that are intermittent. Despite this, our algorithm finds a wider variety of signal morphologies than previous SETI searches focused on narrowband Doppler-drifting signals, while still detecting many such signals. Our pipeline is also less likely than turboSETIto find “candidate” events that clearly contain persistent signals in both ON and OFF scans on visual inspection of spectrograms.
@@
While we use kurtosis as a metric, there are multiple other statistical measures that can be used to evaluate whether a spectrogram differs significantly from Gaussian noise. As one example, the Kullback–Leibler divergence could be calculated instead to determine the deviation of a frequency block’s intensity distribution from the background noise distribution. We leave the implementation of these other approaches to future work.
@@ 6 Conclusion
As part of the Breakthrough Listen program, we searched 2,623 stars, using data from the GBT. We used four receivers, spanning a combined range of 1.10–11.2 GHz, and found no compelling candidates that are not attributable to RFI. We present and implement a novel pipeline, designed to search for signals present in the ON scans of observation cadences and not in the OFF scans. This is a search method less biased towards any specific signal type than many previous SETI searches. We analyze 6,630 unique cadences, representing 3,165 hours of on-sky observation time. We identify $14,168$ frequency blocks 2.86 kHz wide deemed likely to contain a candidate signal. We visually inspected all of these candidates, and determined that all can likely be attributed to RFI. This work highlights the wide variety of RFI that hinders radio SETI searches, and we hope the techniques developed may prove useful in future studies dealing with increasingly heavy RFI environments.
This study constitutes the most comprehensive survey yet for radio evidence of advanced life around nearby stars. We improve on the results of Price et al. (2020) in terms of sensitivity and number of stars. Our novel search approach also makes us sensitive to a larger range of signals than previous surveys. The Drake Figure of Merit of our search is $4.8\times 10^{8}$, the Continuous Waveform Transmittter Rate Figures of Merit of our bands fall between 0.450.850.45-0.85, and we probe down to EIRPs of $2\times 10^{12}$ W for a large numbers of stars (several hundred stars at each band). With respect to each of these metrics, this SETI survey is one of the most sensitive and thorough to date. We add our pipeline to the many algorithms that have been developed to detect technosignatures in radio observations, and hope that this study encourages ensuing SETI surveys to take a more general approach to recovering potential ET signals from complex datasets.
@@ 7 Acknowledgements
The Breakthrough Prize Foundation funds the Breakthrough Initiatives, which manages Breakthrough Listen. The Green Bank Observatory facility is supported by the National Science Foundation, and is operated by Associated Universities, Inc., under a cooperative agreement. CP was funded as a participant in the Berkeley SETI Research Center Research Experience for Undergraduates Site, supported by the National Science Foundation under Grant No. 2244242. The authors thank Joe Bright, Noah Franz, Chenoa Tremblay, and Andrés Luengo for helpful discussions. The authors also thank the anonymous referee and the AAS Journals Statistics Editor for helpful comments which improved the paper.
Appendix A Pointing Direction of Level 2 Blocks and Candidates
Here we plot the RA and declination as well as altitude and azimuth of the sources associated with our final candidates, as well as those associated with our Level 2 Blocks. There appears to be an overdensity of sources located at an azimuth of $210\arcdeg\pm 35\arcdeg$ in L band. This could correspond to a large cluster of geosynchronous satellites as viewed from the GBT. Notably, most of our final candidates come from this cluster.
Figure 16: Telescope pointings of the sources containing our candidate signals.
Figure 17: Telescope pointings of the sources containing our Level 2 Blocks in L Band.
Figure 18: Telescope pointings of the sources containing our Level 2 Blocks in S Band.
Figure 19: Telescope pointings of the sources containing our Level 2 Blocks in C Band.
Figure 20: Telescope pointings of the sources containing our Level 2 Blocks in X Band.
Appendix B Quality Candidates